OpenStudy (anonymous):

Water is pouring into a right circular cone with altitude=12 inches and diameter=4 inches. Which formula below would correctly measure the rate at which the volume is increasing when the altitude is h inches and the altitude is growing at a rate of dh/dt. I know the answer is: dV/dt=1/36(pi)(h)^2(dh/dt) but i need to figure out how to get there.

OpenStudy (amistre64):

Water is pouring into a right circular cone with altitude=12 inches and diameter=4 inches. Which formula below would correctly measure the rate at which the volume is increasing when the altitude is h inches and the altitude is growing at a rate of dh/dt. What is the formula for the volume of a cone?

OpenStudy (anonymous):

V=(1/3)(pi)r^2h

OpenStudy (amistre64):

then we just need to put that thru a derivative to get: V' = (1/3) 2pi r r' dh/dt V' = (2pi r/9) r' dh/dt it looks like they recalculated r in terms of h

OpenStudy (amistre64):

r/h = (4/2)/12 r = h/6 V=(1/3)(pi) r^2 h V=(1/3)(pi) (h/6)^2 h V=(1/3)(pi) h^3/36 V' =(3pi/3)(pi) (h^2/36) dh/dt